OpenStudy (anonymous):

Let f(x) = -4x + 7 and g(x) = 10x - 6. Find f(g(x)). -40x + 24 -40x + 31 -40x + 64 -40x + 70

jimthompson5910 (jim_thompson5910):

Start with the outer function f(x) = -4x + 7

jimthompson5910 (jim_thompson5910):

everywhere you see an x, replace it with g(x) f(x) = -4x + 7 f(g(x)) = -4( g(x) ) + 7

jimthompson5910 (jim_thompson5910):

on the right side, replace g(x) with 10x - 6 since g(x) = 10x - 6 f(g(x)) = -4( g(x) ) + 7 f(g(x)) = -4( 10x - 6 ) + 7 I'll let you finish up. You just need to distribute and combine like terms to simplify.

OpenStudy (anonymous):

ok cool thanks @jim_thompson5910

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

hey is it -40x+31 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

You are correct.